An unknown and unreactive solid has a density of 0.563 g/mL. It is placed in a glass of gasoline. How much of this solid will be above the liquid gasoline if the specific gravity of gasoline is 0.66?
1
Expert's answer
2020-04-21T19:04:21-0400
"\rho_{solid}=563\frac{kg}{m^3},\rho_{gas}=660\frac{kg}{m^3}"

"\rho_{solid}gV=\rho_{gas}gV'"

"\frac{V-V'}{V}=1-\frac{\rho_{solid}}{\rho_{gas}}"

"\frac{V-V'}{V}=1-\frac{563}{660}=0.147\text{ or 14.7\%}"